SPECTROSCOPY OF DIPOLE BOUND ANIONS

\underline{NATHAN HAMMER} and ROBERT COMPTON, Departments of Chemistry and Physics, Buehler Hall, University of Tennessee, Knoxville, TN 37996.

Any molecule with a dipole moment above about 2.5 Debye can form a dipole bound anion. These anions are best produced by charge exchange from atoms in high Rydberg states (Rydberg electron transfer, RET)^{a}. Binding energies (electron affinities, EAs) are estimated from the narrow range of principal quantum number n*_{RET} at which charge exchange occurs or from measurements of the electric field (E_{critical}) required to detach the electron. In one instance (3-methylcyclohexanone) we observe two bound dipole states (3.3 meV and 8.8 meV). An isotope effect in the binding energies of acetone-h (2.92 meV) and acetone-d (2.61 meV) is observed in both n*_{RET} and E_{critical}. Presently, we are studying photodetachment electron spectroscopy of these anions in order to obtain direct values of EA and to access higher excited valence bound anions.

aR. N. Compton and N. I. Hammer ”Multipole-Bound Molecular Anions” in Advances in Gas-Phase Ion Chemistry \underline{4}, edited by N. Adams and L. Babcock, Elsevier Science, 2001.